Solve the following system of simultaneous equations, 6x - 3y = 54 5x + 7y = 7 The solution is ya (Type an integer.)
Answer #1

Answer:

Given:

6x-3y=54                 -------------------equation1

5x+7y=7                  -------------------equation2

Multiply equation 1 by 7 and equation 2 by 3

42x – 21y = 378

15x + 21y = 21

Solving above

57x = 399

x = 7

6x- 3y = 54

6*7 – 3y = 54

3y = -12

Y = -4

Thus value of x= 7 and y = -4